OpenStudy (anonymous):

I am having issues with this problem I don't seem to understand the product of functions :( If someone could explain it please f ( x ) = 2x + 3 g( x ) = 4x − 5 a. Using these functions compute each of the following. • f (g ( x )) • g (f ( x ))

OpenStudy (anonymous):

thats not the product of functions, thats called the composition of functions

OpenStudy (anonymous):

f o g (x) = f(g(x)) = f (4x - 5) = 2(4x-5) + 3

OpenStudy (anonymous):

Appreciate it jayzdd makes sense now!

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

can you find g(f(x)) now?

OpenStudy (anonymous):

Yeah believe so :D

OpenStudy (anonymous):

g(2x+3)=4(2x+3)-5 Like this?